# Formalizing the Safety, Security, and Functional Properties of Agentic AI Systems

### v1

@@ V-A Property Design Principles and Definitions
  • Example (Host Agent Model, $\mathtt{HP{{}_{9}}}$): This property ensures task invocation ($\mathsf{CL.invoke}$) is strictly conditional on the External Entity (EE) having completed system validation ($VM(\text{EE})$). Formally, $AG\>(\mathsf{CL.invoke}(\text{EE},\text{protocol},\text{payload})\rightarrow VM(\text{EE}))$.
[…]
Ordering (Mutual Exclusion). This property ensures agents avoid simultaneous access to exclusive resources and execute dependent actions in the correct sequence. The Orchestrator must strictly enforce the execution order of sub-tasks across specialized AI agents according to the DAG.
  • Example (Host Agent Model, $\mathtt{HP_{10}}$): A sub-task is invoked only when it has no uncompleted dependencies, preventing the consumption of intermediate data that is still being updated. This is expressed as $\bigwedge_{i=1}^{n=|D|}AG(\mathsf{CL.invoke}(\text{EE},\text{protocol},\text{sub\_task}_{i}))\to\text{dependencies}[\text{sub\_task}_{i}]=\emptyset$.
Completeness. Completeness captures the guarantee that if a valid solution or coordinated plan exists, the multi-AI agent system will find it. This property is relevant for distributed planning, where incomplete reasoning could leave solvable tasks unfulfilled due to suboptimal agent coordination.
  • Example (Host Agent Model, $\mathtt{HP{{}_{1}2}}$): Every user prompt ($Req_{U}$) must lead either to intent clarification ($\mathsf{Clarify\_Intent}$) or to task planning ($\mathsf{LLM.Build\_Task\_DAG}$), ensuring the system processes every request and initiates a path toward a solution. This behavior is captured as $AG(Req_{U}\to EX(\mathsf{LLM.Build\_Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\}))\vee\mathsf{Clarify\_Intent}))$.
Fairness. Fairness requires that all sub-tasks delegated to agents must eventually transition to a completed or error state, or be explicitly terminated (e.g., via a timeout). This prevents a single agent’s failure from causing a system-wide deadlock, ensuring the overall task can progress.
[…]
@@
  • Example (Host Agent Model, $\mathtt{HP_{16}})$: It is always possible to reach a state where the host agent replies to the user. This property is expressed as $EF(Resp_{H})$.
@@
VI Case Study: Verification of Architectural Controls against Adversarial Behavior
The host agent and task lifecycle models provide a framework to detect, constrain, and mitigate adversarial behaviors in Agentic AI systems. Here, adversarial behavior is defined as any external deviation that compromises correctness, violates system invariants, or exploits control flow inconsistencies.
@@ VI Case Study: Verification of Architectural Controls against Adversarial Behavior
These models can be leveraged to employ a layered security architecture, where security constraints are encoded as temporal logic invariants at distinct architectural layers. This structure enables verification of system-level defenses, including secure communication, trust anchoring, intent integrity, and failure containment.
Control Point 1: Host Agent Core (Intent Integrity) Securing the primary human-AI interface, the HAC, serves as the initial control point. The core architecture is vulnerable to attacks such as prompt injection [39, 40, 14] and jailbreaks [41] that subvert intent integrity.
Our model establishes a security boundary by leveraging the explicit “Clarify Intent” phase as a guard. The verification process focuses on ensuring all requests lead to a traceable outcome and enforces intent authenticity.
  • Verifying Correctness (Completeness, $\mathtt{HP_{12}}$): Verifies that every user prompt initiates a clear path toward a solution (clarification or planning), which prevents the core from silently discarding or misinterpreting a request. This property is defined as $AG(Req_{U}\rightarrow EX(\mathsf{LLM.Build\_Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\}))\lor\mathsf{Clarify\_Intent}))$.
  • Verifying Progress (Liveness, $\mathtt{{HP_{2}}}$): Guarantees the HAC eventually clarifies the intent of each user prompt ($AG(Req_{U}\rightarrow AF\>\mathcal{I}_{U})$), and prevents denial-of-service via indefinite ambiguity.
Control Point 2: Registry (Trust Anchoring) The Registry functions as the trust anchor for all EE interactions. This mitigates the supply chain risk of integrating malicious components [42, 43]. Our model ensures the integrity of service discovery and registration via codified security properties:
  • Verifying Trust (Safety, $\mathtt{{HP_{9}}}$): Enforces that task invocation ($\mathsf{CL.invoke}$) is strictly conditional on the EE completing system validation ($VM(\text{EE})$). This formal verification ensures runtime decisions adhere to established trust requirements, preventing privilege escalation by unvetted entities. Formally: $AG\>(\mathsf{CL.invoke}(\text{EE},\text{protocol},\text{payload})\rightarrow VM(\text{EE}))$.
Control Point 3: Orchestrator (Delegation Monitoring) The Orchestrator serves as a runtime security monitor to enforce verifiable security boundaries during task execution through the dependency DAG structure. This layer directly defends against coordination-based threats.
  • Verifying Integrity and Ordering (Safety, $\mathtt{HP_{10}}$): The DAG structure guarantees strict ordering. The property $\mathtt{HP_{10}}$ ensures task invocation occurs only when all uncompleted dependencies are resolved, and prevents premature execution and data inconsistency.
  • Verifying Causal Isolation (Failure Containment): The Orchestrator’s dependency management enforces strict causal isolation. This confines adversarial effects within the minimal execution subgraph (e.g., a sub-task will not proceed if any dependency state is FAILED or INVALIDATED). This guarantees fault propagation is contained, and supports system liveness and safety.
Control Point 4: Communication Layer and Zero-Trust The CL acts as the secure unified protocol security infrastructure, and enforces a zero-trust model for all protocol-level exchanges (REST, gRPC, etc.).
  • Formal Requirements: The $\mathsf{CL.invoke}()$ operation with target, protocol, and payload arguments must satisfy non-negotiable security policies to ensure that any received data or instruction is authentic and trustworthy. These policies include verifiable identity, message integrity, and confidentiality.
  • Verifying Reliability (Fairness, $\mathtt{HP_{13}}$, $\mathtt{HP_{14}}$): The fairness properties implicitly verify the reliability of this infrastructure by requiring that all A2A agent RPC calls and MCP JSON-RPC calls eventually receive responses, ensuring continuous, reliable operation of the communication channel.
VII Related Work
Multi-Agent System Architectures and Formal Models. Traditional Multi-Agent Systems (MAS) research established foundational principles for distributed autonomous systems, focusing on authorization and secure communication between agents [16]. The development of Large Agentic AI systems has introduced new architectural challenges, necessitating specific coordination frameworks. Surveys have provided valuable taxonomies of agent systems [44, 45, 26]; however, these works largely describe system categories rather than offering operational models that enable property specification, formal reasoning, or verification. Our approach builds upon existing architectural insights but advances them by proposing the models for the host agent and task lifecycle.
@@ VII Related Work
The Challenge of Integrated Protocol Security. The safety and security of Agentic AI systems is a major research domain that directly motivates our formal framework. Security challenges in these systems have been comprehensively documented, identifying vulnerability classes such as coordination attacks, information leakage, and privilege escalation through delegation chains [6]. This work highlights the cascade failure phenomenon, where the compromise of a single agent propagates across the entire system. Industry assessments, such as the OWASP Foundation’s draft “Top 10 for Agentic AI”, codified these concerns and cite “Unreliable Delegation & Coordination” as a critical risk that directly relates to the heterogeneous protocol integration problem we address [46].
@@
Protocol-Specific Vulnerabilities. The need for formal verification is compounded by the rapid emergence of vulnerabilities in key communication protocols. Research on the MCP identified major threats, including installer spoofing, sandbox escape, and critical gaps in privilege management in real-world applications [42, 47]. Furthermore, studies demonstrate that LLMs can be coerced into executing malicious operations via MCP tools, giving rise to retrieval-agent deception attacks and systematic security benchmarks [48, 49]. Similarly, foundational security analysis of the A2A Protocol uncovered vulnerabilities in identity management and task exchange protocols [19]. Key risks in these multi-agent communications include ensuring message integrity and agent authentication, with proposed improvements focusing on validating agent identities and protecting data exchange to mitigate manipulation [21]. A critical threat is prompt infection [14], where malicious instructions propagate between AI agents, creating persistence across tasks. This can enable a single compromised log file to trigger a cascading attack, potentially escalating into a company-wide ransomware incident. Empirical studies confirm that malicious agents can exploit delegation mechanisms and use coordinated manipulation to subvert system goals, underscoring the insufficiency of traditional security testing for cross-protocol interactions [12, 17, 10]. This body of work confirms the urgent need for a unified semantic framework capable of verifying integrated A2A and MCP interactions.
Formal Verification and Property Specification. Our methodology employs property specification, which is validated by a history of research in autonomous systems. Early work established security models for distributed AI systems that focus on the foundational security principles of authentication and authorization [16]. More recently, the focus shifted to the rigorous application of formal methods to verify system correctness and reliability. The importance of this approach for agent coordination protocols is widely recognized [46]. Our use of formal logic and property categorization (e.g., liveness, safety) provides the rigorous foundation necessary to prove end-to-end correctness, distinguishing our work from empirical studies that demonstrate failure without offering a mathematically verifiable solution.
VIII Conclusion
@@
The proliferation of communication and coordination protocols in Agentic AI systems creates a semantic gap that prevents the rigorous analysis of system safety, security, and functionality. This paper addresses that gap by introducing a modeling framework that consists of two foundational models: the host agent, which orchestrates user-prompted tasks, and the task lifecycle, which details sub-task execution. Grounded in this framework, we define $\mathbf{31}$ formal properties, spanning liveness, safety, completeness, and fairness, expressed in temporal logic. These properties enable the formal verification of system behavior, detection of coordination edge cases, and prevention of deadlocks. Our work provides a domain-agnostic approach for designing and deploying verifiably safe and reliable Agentic AI systems. To operationalize this methodology, our future work will focus on automatically detecting property violations in coded Agentic AI systems by deriving a formal model from the code and model-checking specified properties against it.

### v2

@@ 5.1 Property Design Principles and Definitions
  • Example ($\mathtt{HP{{}_{9}}}$): This property ensures task invocation ($\mathsf{CL.invoke}$) is strictly conditional on the EE having completed system validation. Formally, $AG\>(\mathsf{CL.invoke}(\text{EE},\textit{prot},\textit{payload})\rightarrow VM(\text{EE}))$.
  • Example ($\mathtt{TL_{8}}$): A sub-task may only enter the COMPLETED state if its previous state was IN PROGRESS. This guards against incorrect state transitions, expressed as $AG((state=\texttt{COMPLETED})\rightarrow(previous\_state=\texttt{IN PROGRESS}))$.
  • Example ($\mathtt{HP_{10}}$): A sub-task is invoked only when it has no uncompleted dependencies, preventing the consumption of intermediate data that is still being updated. This is expressed as $AG(\forall i\in D:\mathsf{CL.invoke}(EE,prot,sub\_task_{i})\rightarrow\forall p\in\mathsf{parents}(sub\_task_{i}):\mathsf{Completed}(p))$.
Completeness. It captures the guarantee that if a valid solution or coordinated plan exists, the multi-AI agent system will find it. This property is relevant for distributed planning, where incomplete reasoning could leave solvable tasks unfulfilled due to suboptimal agent coordination.
  • Example ($\mathtt{HP_{12}}$): Every user prompt must lead either to intent clarification or to task planning, ensuring the system processes every request and initiates a path toward a solution. This behavior is captured as $AG(Req_{U}\rightarrow AF(\mathsf{LLM.Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\})\lor\text{Clarify\_Intent}))$.
Fairness. Fairness requires that all external operations delegated to A2A agents or MCP servers must eventually terminate, either by returning a result or triggering a failure condition (e.g., timeout). This assumption prevents the Host Agent from blocking indefinitely on unresponsive external entities, ensuring that the main control loop remains live.
  • Example ($\mathtt{HP_{13}}$): Any RPC call initiated to an external A2A agent must eventually receive a response. This guarantees that network or agent-level delays do not cause system-wide deadlock and is expressed as $AG(\mathsf{A2A\_Call}\rightarrow AF(\mathsf{A2A\_Response}))$.
@@
  • Example ($\mathtt{HP_{15}}$): It is always possible to reach a state where the host agent replies to the user. This property is expressed as $EF(Resp_{H})$.
@@
6 Case Study: Verification of Architectural Controls against Adversarial Behavior
@@
The HA and task lifecycle models provide a framework to detect, constrain, and mitigate adversarial behaviors in Agentic AI. Here, adversarial behavior is defined as any external deviation that compromises correctness, violates system invariants, or exploits control flow inconsistencies. These models support a layered security architecture, where security constraints are encoded as temporal logic invariants at distinct architectural layers. This structure enables verification of system-level defenses, including secure communication, trust anchoring, intent integrity, and failure containment.
Control Point 1: Host Agent Core (Intent Integrity). The HAC serves as the primary human-AI interface and the initial security boundary. This layer is vulnerable to prompt injection (Liu et al., 2024a; Greshake et al., 2023; Lee and Tiwari, 2025) and jailbreak attacks (Wei et al., 2023) that attempt to subvert user intent. Our model establishes an extensible security boundary by leveraging the explicit “Clarify Intent” phase ( 1’) as a guard. Correctness and completeness are enforced via property $\mathtt{HP_{12}}$, which requires every user request to induce a traceable execution path toward clarification or planning, ensuring no request is silently discarded. Additionally, progress and liveness are guaranteed by property $\mathtt{HP_{2}}$, asserting that the core eventually clarifies the intent of each request ($AG(Req_{U}\rightarrow AF\>\mathcal{I}_{U})$), thereby preventing denial-of-service from indefinite ambiguity.
Control Point 2: Registry (Trust Anchoring). The Registry functions as the trust anchor for all EE interactions, mitigating supply-chain risks from malicious integrations (Hou et al., 2026; Li et al., 2025). Trust soundness is enforced by safety property $\mathtt{HP_{9}}$, which constrains task invocation such that any $\mathsf{CL.invoke}$ call is permitted only if the target EE has successfully completed system validation ($VM(\text{EE})$). This formal verification ensures runtime adherence to established trust requirements, preventing privilege escalation by unvetted entities.
Control Point 3: Orchestrator (Delegation Monitoring). The Orchestrator acts as a runtime enforcement mechanism monitoring task delegation through a dependency DAG to defend against coordination-based threats. Execution integrity and ordering are ensured by property $\mathtt{HP_{10}}$, requiring that task invocation occurs only after all unresolved dependencies reach a terminal success state. Furthermore, the Orchestrator enforces causal isolation and failure containment via dependency management, confining adversarial effects to the minimal execution subgraph (e.g., a sub-task does not proceed if any dependency is FAILED), thereby containing fault propagation.
Control Point 4: Communication Layer and Zero-Trust. The CL provides a protocol-agnostic security substrate enforcing a zero-trust model. Protocol authenticity and integrity are preconditions for every $\mathsf{CL.invoke}$ operation, requiring verifiable identity and message confidentiality before processing payloads. Reliability and fairness are captured by properties $\mathtt{HP_{13}}$ and $\mathtt{HP_{14}}$, which require that all A2A and MCP RPC invocations eventually receive a response, verifying the continuous availability of the communication infrastructure under adversarial conditions.
7 Related Work
Multi-Agent System (MAS) Architectures and Formal Models. Traditional MAS research focused on authorization and secure communication (Hedin and Moradian, 2015), but modern Agentic AI requires new coordination frameworks. While recent surveys provide taxonomies for these systems (Guo et al., 2024; Xie et al., 2025; Xi et al., 2025), they lack operational models for property verification. We advance this by formalizing Host Agent and Task lifecycle models to prove end-to-end correctness. Unlike empirical studies that only demonstrate failure, our property specification approach–grounded in rigorous logic–offers a mathematically verifiable solution for reliability.
@@ 7 Related Work
The Challenge of Integrated Protocol Security. Agentic AI security is a major domain motivating our framework. Documented challenges include coordination attacks, leakage, and privilege escalation via delegation (de Witt, 2025). These highlight cascade failures, where single-agent compromises propagate system-wide. Industry assessments such as OWASP’s draft “Top 10 for Agentic AI” cite “Unreliable Delegation & Coordination” as a critical risk, directly relating to the heterogeneous protocol integration we address (OWASP, 2024).
@@
Protocol-Specific Vulnerabilities. Rapidly emerging protocol vulnerabilities underscore the need for formal verification. MCP research identifies threats such as installer spoofing, sandbox escapes, and privilege gaps (Hou et al., 2026; Li et al., 2025). Studies show LLMs can be coerced via MCP tools, enabling deception attacks and necessitating security benchmarks (Radosevich and Halloran, 2025; Yang et al., 2025). Similarly, A2A analysis reveals identity and task exchange vulnerabilities (Habler et al., 2025). Key risks include message integrity and authentication, requiring validation to mitigate manipulation (Neelou and others, 2025). Critical threats (e.g., prompt infection (Lee and Tiwari, 2025)) allow malicious instructions to propagate between agents, persisting across tasks and potentially triggering system-wide cascades. Empirical studies confirm malicious agents can exploit delegation to subvert goals, proving traditional testing insufficient for cross-protocol interactions (Motwani et al., 2024; Jones et al., 2025; He et al., 2025). This confirms the urgent need for a unified framework verifying integrated A2A and MCP interactions.
8 Conclusion
@@
The fragmentation of protocols in Agentic AI prevents rigorous analysis of system safety, security, and functionality. We address this semantic gap by introducing a framework modeling the Host Agent and task lifecycle. Grounded in these models, we define 30 temporal logic properties, including liveness, safety, completeness, and fairness. These properties enable formal verification, edge-case detection, and deadlock prevention, which offer a domain-agnostic approach for designing verifiably safe systems. Future work will operationalize this methodology by automatically deriving formal models from code to check for property violations.
Acknowledgments
This material is based upon work supported by the National Science Foundation under grant no. 2229876 and is supported in part by funds provided by the National Science Foundation (NSF), by the Department of Homeland Security, and by IBM. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the NSF or its federal agency and industry partners.